Help with 2 please

A 50 liter tank contains a water-vapor mixture at
Answer #1

final percentage = 25+1.89 = 26.89 litres

thus % = (26.89/50) X 100 =53.8 %

5 atonわ知|b| 2257 22s? Heat for. Ik